OpenStudy (anonymous):

The length of a rectangle is 2 more than twice the width. Find the dimensions of the rectangle if the perimeter is 76. If x represents the width of the rectangle, then which expression represents the perimeter? would the answer be: 4x + 4?

OpenStudy (matt101):

If x is the width, the length is 2x+2 (two more than twice the width). if the perimeter is 76 units, then: 2L + 2W = 76 2(2x+2) + 2x = 76 6x+4=76 x = 12 The width is 12 units and the length is 2(12)+2=26 units CHECK: 2(12)+2(26)=76 units :)